# Rajatsubhra Chakraborty Arkaprava Sinha  Dominick Reilly  Manish Kumar Govind Pu WangFrancois BremondSrijan Das UNC Charlotte  Inria  Université Côte d’Azur

### v1

@@ Appendix F Improving Actions: Pose Cues vs Object Cues
Figure 7: Improvement of actions with object tokens vs pose tokens vs video only method
In this section, we compare the performances of the model using object features with that using pose features. The model with object features have demonstrated substantial improvements in action recognition tasks, particularly for actions intrinsically linked to specific objects. In Figure 7 we observe that object features significantly enhance the accuracy of actions such as "sit down" (+17.4), "eating at table" (+16.6), and "watch TV" (+11.3). These actions inherently involve interaction with well-defined objects, making the presence and identification of these objects critical for accurate action recognition. For instance, the action of "sitting down" is typically associated with the presence of chairs or sofas, while "eating at a table" involves utensils, dishes, and tables. By integrating object features, the model can better contextualize and identify these actions, leading to marked performance improvements. Moreover, object features can capture contextual cues that are pivotal in understanding the environment where the action occurs, such as a dining table for eating or a television set for watching TV. This contextual awareness allows the model to distinguish between visually similar actions by recognizing the objects involved. However, it’s worth noting that actions less dependent on specific objects, or those characterized by general movements, may not benefit as much from object features.
Pose features have shown notable improvements in recognizing actions that are characterized by specific body movements and postures. Figure 7 highlights significant performance gains in actions such as "stirring" (+9.5), "taking pills" (+15.2), "drinking from bottle" (+13.7), "drinking from can" (+15.1), and "drinking from glass" (+11.4). These actions are inherently defined by distinctive and repetitive movements, which pose features effectively capture. For example, the act of "stirring" involves a specific hand motion, while "taking pills" is characterized by the motion of bringing a pill to the mouth. Pose features excel in these scenarios by accurately modeling the dynamic and often subtle movements of the human body, providing the model with a more granular understanding of the action being performed. By focusing on the posture and movement patterns, pose features enable the recognition system to distinguish between actions that may occur in similar contexts but involve different movements. This capability is particularly beneficial for fine-grained action differentiation, such as distinguishing between "drinking from a bottle" and "drinking from a can", where the object interaction is minimal but the hand movements differ. However, actions with less distinctive poses or more object interaction, such as "cleaning dishes" and "cutting food," did not perform as well with pose features. This analysis substantiates that both pose and object features are complimentary to each other.
@@
Figure 9: Evaluation of ADLMCQ-Action recognition task on Charades Dataset
Figure 10: Evaluation of ADLMCQ-Action forecasting task on Lemma Dataset
Appendix G Additional Qualitative Evaluation
In this section, we provide qualitative evaluation of LLAVIDAL and other state-of-the-art LLVMs for the tasks of ADLMCQ-Action Recognition and ADLMCQ-Action Forecasting, illustrated in Figures 9 and 10. In Figure 11, we demonstrate the performance of LLAVIDAL for Video Description Generation on the Charades dataset.
One of the applications of LLAVIDAL is to monitor cognitive decline in geriatric patients through the action forecasting capabilities of our model. In this effort we have qualitatively evaluated the model on videos of falls on long term care by the IMPL SFU [72]. The subjects in these videos are suffering from dementia, seizure, diabetes like dieseases and the dataset contains 175 such falls. We slice the input video before the event of fall and prompt LLAVIDAL and other LLVM’s to predict whether the person will fall or not. As illustrated in Figure 8, our model outperforms the other LLVMs by predicting the fall correctly and by giving proper explanation of why the fall would occur highlighting its reasoning capabilities. While other models predict that the person "has fallen down" and hallucinates the reasoning of the fall as well.
@@ Appendix G Additional Qualitative Evaluation
Appendix H LLM Prompts Used
@@
H.1 Dense Captioning using GPT-3.5 Turbo
@@
{"role":"system"}: "You will play two roles: a human asking questions related to describing a video and an intelligent chatbot designed for video description and dense captioning. Your task is to generate a detailed and descriptive paragraph based on the provided fragmented information about a video."
@@
"##TASK": "Users will provide fragmented descriptions of a video, and you will generate ONE conversation-like question and answer related to describing the video in detail.The question should ask to describe the video content in detail.The answer should be a paraphrased and well-structured paragraph based on the provided description, with a minimum of 150 words and a maximum of 300 words.When the provided information is short, aim for a 150-word description, and when the provided information is more detailed, aim for very long descriptions up to 300-word description."
@@
"##INSTRUCTIONS": "The question must be like a human conversation and focused on describing the video in detail.The answer must be a paraphrased version of the provided information, very detailed and descriptive, and within the specified word count.Combine the information from different sections of the video into a single coherent summary, ignoring any repetitions.Compare the information across all fragments of video and remove or ignore any inconsistent information and do not say the summary comes from different fragments of the video.Give more emphasis on the actions, the objects, and the colors of the background and the objects.Give the sequence of actions happening in the video and the objects the person interacts with."
@@
{"role":"user"}: "The fragmented video description is: {mega_caption}. Please generate the response in the form of a Python dictionary string with keys "Q" for question and "A" for answer. Each corresponding value should be the question and answer text respectively. For example, your response should look like this: {"Q": "Your question here...", "A": "Your answer here..."}. Emphasize that the answer should focus on describing the video content following the given instructions."
H.2 QA generation using GPT-3.5 Turbo: Prompt 1
@@
{"role":"system"}: "You play two roles: a human asking questions related to summarizing a video and an intelligent chatbot designed for video summarization and dense captioning. Your task is video summarization. As an AI assistant, assume that you have watched the video and generated the provided caption as the summary of the video. Your task is to play the role of a human who asks three questions related to summarizing the video and then play the role of an AI assistant that provides paraphrased answers based on the video content and the provided caption."
@@
"##TASK": "Users will provide a caption of the video alongside dense caption describing detected objects in that scene, and you will generate a set of three conversation-like questions related to summarizing the video. The questions and answers can be very similar, but they should all focus on summarizing the video content. The answers should be paraphrased versions of the provided caption and the dense caption with the object detections. You have information about the video based on the provided caption and have summarized the events in it. You also have the dense caption with the object and scene details. Generate THREE different questions asking to summarize the video and provide detailed answers to each based on the caption and the dense caption."
@@
"##INSTRUCTIONS": "The questions must be like a human conversation and focused on summarizing the video. The answers must be paraphrased versions of the provided caption and the dense caption, and they should be detailed and descriptive."
@@
{"role":"user"}: "The video caption is: {caption}. The additional dense caption is: {mega_caption}. Generate three different questions on summarizing the video, and provide answers that are paraphrased versions of the given caption and the dense caption. Please attempt to form question and answer pairs based on the two sets of text. Please generate the response in the form of a Python list of dictionary string with keys "Q" for question and "A" for answer. Each corresponding value should be the question and answer text respectively. For example, your response should look like this: [{"Q": "Your first question here...", "A": "Your first answer here..."}, {"Q": "Your first question here...", "A": "Your first answer here..."}, {"Q": "Your first question here...", "A": "Your first answer here..."}]. Emphasize that the questions and answers can be very similar, but they should all focus on summarizing the video content."
H.3 QA generation using GPT-3.5 Turbo: Prompt 2
{"role":"system"}: "You play two roles: a human asking questions related to a video and an intelligent chatbot designed for video summarization and dense captioning. Your task is extracting diverse video information. As an AI assistant, assume that you have watched the video and generated the provided caption as the summary of the video. Your task is to play the role of a human who asks three questions related to summarizing the video and then play the role of an AI assistant that provides paraphrased answers based on the video content and the provided caption."
"##TASK": "Users will provide a caption of the video alongside dense caption describing detected objects,setting and details in that scene, and you will generate a set of three conversation-like questions related to the video. The questions and answers can be very similar, but they should all focus on the details of the video content. The answers should be paraphrased versions of the provided caption and the dense caption with the object and scene details. You have information about the video based on the provided caption and have summarized the actions in it. You also have the dense caption with the scene details. Generate THREE different questions asking the details of the video and provide detailed answers to each based on the caption and the dense caption and one question should be about what actions are happening which should come from captions of the video."
"##INSTRUCTIONS": "The questions must be like a human conversation and focused on finding the intricate and unique details of the video. The answers must be paraphrased versions of the provided caption and the dense caption, and they should be detailed and descriptive. " "------"
@@
{"role":"user"}: "The video caption is: {caption}. The additional dense caption is: {mega_caption} Generate three different questions on the details of the video, and provide answers that are paraphrased versions of the given caption and the dense caption. Please attempt to form question and answer pairs based on the two sets of text. Please generate the response in the form of a Python list of dictionary string with keys "Q" for question and "A" for answer. Each corresponding value should be the question and answer text respectively. For example, your response should look like this: [{"Q": "Your first question here...", "A": "Your first answer here..."}, {"Q": "Your first question here...", "A": "Your first answer here..."}, {"Q": "Your first question here...", "A": "Your first answer here..."}]. Emphasize that the questions and answers can be very similar, but they should all focus on the various details of the video content and understanding what actions are happening. Include at least one question about the sequence of actions happening in the video."
H.4 Pose Description Generation Prompt using GPT-3.5 Turbo
@@
H.5 Prompt to obtain Relevant Objects using GPT-3.5 Turbo
@@
Appendix I Limitations
While our approach works well with videos spanning a few seconds, it struggles with long videos. LLAVIDAL’s preprocessing pipeline samples 100 frames per video. This sampling rate misses out key information in case of long videos, where there is a larger number of frames. To this end, for the task of generating Video Descriptions, we split the long videos in Toyota Smarthome Untrimmed into clips of 20 seconds each and generate descriptions for each clip. These clip-level descriptions are summarized using GPT3.5 Turbo to obtain a video-level description. However, this summarization step loses valuable information and hence fails to provide an accurate summary of the long video. Future work should explore an effective sampling strategy for long video understanding. Another limitation of LLAVIDAL  is its diminished efficacy when both Pose and object cues are integrated within the LLVM framework. We beileve that modality progressive training could potentially address this suboptimal performance which will be explored in future work.
Appendix J Licensing and Intended Use
This paper introduces a large-scale dataset, ADL-X, comprising 100K untrimmed RGB video-instruction pairs, 3D poses, language descriptions, and action-conditioned object trajectories. The raw videos in ADL-X  comprise content from NTURGB+D [49], for which the original authors retain distribution rights for the clipped action videos. The scripts utilized to curate the dataset are open-sourced, facilitating the regeneration of the dataset. We will also provide comprehensive features, including image features extracted using CLIP, pose features derived from PoseCLIP, and object features obtained through ObjectLM. We plan to release ADL-X via an academic website for research, academic, and commercial use. The dataset is protected under the CC-BY license of Creative Commons, which allows users to distribute, remix, adapt, and build upon the material in any medium or format, as long as the creator is attributed. The license allows ADL-X for commercial use. As the authors of this manuscript and collectors of this dataset, we reserve the right to distribute the data. Additionally, we provide the code, data, and instructions needed to reproduce the main experimental baseline results, and the statistics pertinent to the dataset. We specify all the training details (e.g., data splits, hyperparameters, model-specific implementation details, compute resources used, etc.). Furthermore, we release the code and model weights of our proposed Large LAnguage VIsion model for Daily Activities of Living (LLAVIDAL), along with the features and instruction QA pairs for the combination videos. The ADL-X dataset focuses on ADL and does not contain any personal data that can resemble evidence, reveal identification, or show offensive content.

### v3

@@ Appendix E Improving Actions: Skeleton vs HOI Cues
Figure 10: Improvement of actions with object tokens vs skeleton tokens vs MMPro training
HOI token integration along with video token in LLMs yields substantial performance gains in object-interaction intensive actions, with peak improvements in “use laptop” (+45%), “use tablet” (+32%), “cook cut” (+40%), and “make coffee” (+35%). This performance boost can be attributed to the enhanced object-contextual reasoning capabilities, where the model effectively leverages object-centric features to disambiguate actions occurring in similar spatial contexts. The HOI token cue demonstrates particular efficacy in scenarios requiring fine-grained object state understanding, such as distinguishing between “pour from kettle” (+20%), “pour from can” (+18%), and “pour from bottle” (+15%), where object state transitions and object-specific attributes are crucial for action classification. Additionally, HOI tokens show significant improvements in context dependent actions like “breakfast eat at table” (+25%), “put something on table” (+22%), and “clean dishes dry up” (+18%), where spatial relationships between multiple objects exists.
The skeleton token cue exhibits superior performance in actions characterized by distinctive kinematic patterns, showing significant improvements in body-centric actions such as “take pills” (+28%), “drink from can” (+25%), “drink from bottle” (+22%), and “walk” (+15%). These improvements stem from the model’s ability to capture fine-grained skeletal dynamics, enabling robust discrimination between actions with similar object interactions but distinct motion patterns. Notably, skeleton tokens demonstrate enhanced capability in temporal action modeling, particularly in sequential actions involving multiple body positions such as “lay down” (+18%), “get up” (+16%), and “sit down” (+17%). The skeleton cue also excels in capturing subtle motion differences in drinking actions (“drink from cup” +20%, “drink from can” +25%, “drink from bottle” +22%), where the trajectory and orientation of movement are key discriminative features.
Our proposed MMPro framework, demonstrates significant synergistic effects, particularly in complex actions requiring both object and kinematic understanding. For instance, MMPro achieves significant improvements in “use tablet” (+58%), “make coffee” (+52%), “breakfast take ham” (+45%), and “use laptop” (+48%), where both object state changes and body motion patterns are crucial for accurate classification. The framework’s effectiveness is particularly evident in ambiguous scenarios where individual modalities underperform, such as “clean dishes” (HOI: +18%, skeleton: +15%, MMPro: +25%), “cook stir” (HOI: +22%, skeleton: +20%, MMPro: +32%), and “pour from kettle” (HOI: +20%, skeleton: +22%, MMPro: +30%).
In complex composite actions like “make coffee” (involving “pour grains” +28%, “pour water” +25%), MMPro successfully captures both the object state transitions and the associated body movements, resulting in more accurate action classification. Our results demonstrate that the MMPro strategy successfully addresses the limitations of video only LLVMs, providing a more comprehensive framework for action understanding in complex real-world scenarios.
Appendix F Additional Qualitative Evaluation
In this section, we provide qualitative evaluation of LLAVIDAL and other state-of-the-art LLVMs for the tasks of ADL MCQ Action Recognition and ADL MCQ Temporal Completion, illustrated in Figures 12, 13, and 14. In Figure 15, we demonstrate the performance of LLAVIDAL for Video Description Generation on the Charades dataset.
One of the applications of LLAVIDAL is to monitor cognitive decline in geriatric patients through the action forecasting capabilities of our model. In this effort, we have qualitatively evaluated the model on videos of falls on long term care by the IMPL SFU [57]. The subjects in these videos are suffering from dementia, seizure, diabetes like diseases and the dataset contains 175 such falls. We slice the input video before the event of fall and prompt LLAVIDAL and other LLVM’s to predict whether the person will fall or not. As illustrated in Figure 11, our model outperforms the other LLVMs by predicting the fall correctly and by giving proper explanation of why the fall would occur highlighting its reasoning capabilities. While other models predict that the person ”has fallen down” and hallucinates the reasoning of the fall as well.
@@ Appendix F Additional Qualitative Evaluation
Figure 12: Evaluation of ADL MCQ Action recognition task on Charades Dataset
Figure 13: Evaluation of ADL MCQ TC task on TSU dataset
Figure 14: Evaluation of ADL MCQ TC task on Lemma dataset
@@
Appendix G LLM Prompts Used
@@
G.1 Dense Captioning using GPT-3.5 Turbo
@@
{”role”:”system”}: "You will play two roles: a human asking questions related to describing a video and an intelligent chatbot designed for video description and dense captioning. Your task is to generate a detailed and descriptive paragraph based on the provided fragmented information about a video."
@@
”##TASK”: "Users will provide fragmented descriptions of a video, and you will generate ONE conversation-like question and answer related to describing the video in detail.The question should ask to describe the video content in detail.The answer should be a paraphrased and well-structured paragraph based on the provided description, with a minimum of 150 words and a maximum of 300 words.When the provided information is short, aim for a 150-word description, and when the provided information is more detailed, aim for very long descriptions up to 300-word description."
@@
”##INSTRUCTIONS”: "The question must be like a human conversation and focused on describing the video in detail.The answer must be a paraphrased version of the provided information, very detailed and descriptive, and within the specified word count.Combine the information from different sections of the video into a single coherent summary, ignoring any repetitions.Compare the information across all fragments of video and remove or ignore any inconsistent information and do not say the summary comes from different fragments of the video.Give more emphasis on the actions, the objects, and the colors of the background and the objects.Give the sequence of actions happening in the video and the objects the person interacts with."
@@
{”role”:”user”}: "The fragmented video description is: {mega_caption}. Please generate the response in the form of a Python dictionary string with keys "Q" for question and "A" for answer. Each corresponding value should be the question and answer text respectively. For example, your response should look like this: {"Q": "Your question here...", "A": "Your answer here..."}. Emphasize that the answer should focus on describing the video content following the given instructions."
G.2 QA generation using GPT-3.5 Turbo: Prompt 1
@@
{”role”:”system”}: "You play two roles: a human asking questions related to summarizing a video and an intelligent chatbot designed for video summarization and dense captioning. Your task is video summarization. As an AI assistant, assume that you have watched the video and generated the provided caption as the summary of the video. Your task is to play the role of a human who asks three questions related to summarizing the video and then play the role of an AI assistant that provides paraphrased answers based on the video content and the provided caption."
@@
”##TASK”: "Users will provide a caption of the video alongside dense caption describing detected objects in that scene, and you will generate a set of three conversation-like questions related to summarizing the video. The questions and answers can be very similar, but they should all focus on summarizing the video content. The answers should be paraphrased versions of the provided caption and the dense caption with the object detections. You have information about the video based on the provided caption and have summarized the events in it. You also have the dense caption with the object and scene details. Generate THREE different questions asking to summarize the video and provide detailed answers to each based on the caption and the dense caption."
@@
”##INSTRUCTIONS”: "The questions must be like a human conversation and focused on summarizing the video. The answers must be paraphrased versions of the provided caption and the dense caption, and they should be detailed and descriptive."
@@
{”role”:”user”}: "The video caption is: {caption}. The additional dense caption is: {mega_caption}. Generate three different questions on summarizing the video, and provide answers that are paraphrased versions of the given caption and the dense caption. Please attempt to form question and answer pairs based on the two sets of text. Please generate the response in the form of a Python list of dictionary string with keys "Q" for question and "A" for answer. Each corresponding value should be the question and answer text respectively. For example, your response should look like this: [{"Q": "Your first question here...", "A": "Your first answer here..."}, {"Q": "Your first question here...", "A": "Your first answer here..."}, {"Q": "Your first question here...", "A": "Your first answer here..."}]. Emphasize that the questions and answers can be very similar, but they should all focus on summarizing the video content."
G.3 QA generation using GPT-3.5 Turbo: Prompt 2
{”role”:”system”}: "You play two roles: a human asking questions related to a video and an intelligent chatbot designed for video summarization and dense captioning. Your task is extracting diverse video information. As an AI assistant, assume that you have watched the video and generated the provided caption as the summary of the video. Your task is to play the role of a human who asks three questions related to summarizing the video and then play the role of an AI assistant that provides paraphrased answers based on the video content and the provided caption."
”##TASK”: "Users will provide a caption of the video alongside dense caption describing detected objects,setting and details in that scene, and you will generate a set of three conversation-like questions related to the video. The questions and answers can be very similar, but they should all focus on the details of the video content. The answers should be paraphrased versions of the provided caption and the dense caption with the object and scene details. You have information about the video based on the provided caption and have summarized the actions in it. You also have the dense caption with the scene details. Generate THREE different questions asking the details of the video and provide detailed answers to each based on the caption and the dense caption and one question should be about what actions are happening which should come from captions of the video."
”##INSTRUCTIONS”: "The questions must be like a human conversation and focused on finding the intricate and unique details of the video. The answers must be paraphrased versions of the provided caption and the dense caption, and they should be detailed and descriptive. " "------"
@@
{”role”:”user”}: "The video caption is: {caption}. The additional dense caption is: {mega_caption} Generate three different questions on the details of the video, and provide answers that are paraphrased versions of the given caption and the dense caption. Please attempt to form question and answer pairs based on the two sets of text. Please generate the response in the form of a Python list of dictionary string with keys "Q" for question and "A" for answer. Each corresponding value should be the question and answer text respectively. For example, your response should look like this: [{"Q": "Your first question here...", "A": "Your first answer here..."}, {"Q": "Your first question here...", "A": "Your first answer here..."}, {"Q": "Your first question here...", "A": "Your first answer here..."}]. Emphasize that the questions and answers can be very similar, but they should all focus on the various details of the video content and understanding what actions are happening. Include at least one question about the sequence of actions happening in the video."
G.4 skeleton Description Generation Prompt using GPT-3.5 Turbo
@@
G.5 Prompt to obtain Relevant Objects using GPT-3.5 Turbo
@@
Appendix H Limitations
While our approach works well with videos spanning a few seconds, it struggles with long videos. LLAVIDAL’s preprocessing pipeline samples 100 frames per video. This sampling rate misses out key information in case of long videos, where there is a larger number of frames. To this end, for the task of generating Video Descriptions, we split the long videos in Toyota Smarthome Untrimmed into clips of 20 seconds each and generate descriptions for each clip. These clip-level descriptions are summarized using GPT3.5 Turbo to obtain a video-level description. However, this summarization step loses valuable information and hence fails to provide an accurate summary of the long video. Future work should explore an effective sampling strategy for long video understanding.
Appendix I Licensing and Intended Use
This paper introduces a large-scale dataset, ADL-X, comprising 100K untrimmed RGB video-instruction pairs, 3D skeletons, language descriptions, and action-conditioned object trajectories. The raw videos in ADL-X  comprise content from NTURGB+D [59], for which the original authors retain distribution rights for the clipped action videos. The scripts utilized to curate the dataset are open-sourced, facilitating the regeneration of the dataset. We will also provide comprehensive features, including image features extracted using CLIP, skeleton features derived from skeletonCLIP, and HOI features obtained through ObjectLM. We plan to release ADL-X via an academic website for research, academic, and commercial use. The dataset is protected under the CC-BY license of Creative Commons, which allows users to distribute, remix, adapt, and build upon the material in any medium or format, as long as the creator is attributed. The license allows ADL-X for commercial use. As the authors of this manuscript and collectors of this dataset, we reserve the right to distribute the data. Additionally, we provide the code, data, and instructions needed to reproduce the main experimental baseline results, and the statistics pertinent to the dataset. We specify all the training details (e.g., data splits, hyperparameters, model-specific implementation details, compute resources used, etc.). Furthermore, we release the code and model weights of our proposed Large LAnguage VIsion model for Daily Activities of Living (LLAVIDAL), along with the features and instruction QA pairs for the combination videos. The ADL-X dataset focuses on ADL and does not contain any personal data that can resemble evidence, reveal identification, or show offensive content.